OpenStudy (anonymous):

Find all the solutions Csc^2 x- csc x-2=0

OpenStudy (anonymous):

Solve y^2 - y -2 = 0 y = [1+- sqrt(1+8)]/2 = [1+-3]/2 = -1, 2 csc x = -1, 2 1/sin x = -1, 2 sin x = -1, 1/2 x = etc.

OpenStudy (anonymous):

Thanks. Glad to have helped.

OpenStudy (anonymous):

:)